# Searching for axion dark matter with magnetic resonance force microscopy

### v1

@@ Appendix A Appendix A: Dark Matter Couplings
Dark Photon Kinetic Mixing.– The dark photon ${A^{\prime}}^{\mu}$ can kinetically mix with with the Standard Model photon $A^{\mu}$, leading to the following interaction operator in the interaction basis fedderke2021earth; kalia2024ultralight
@@
where $\epsilon$ is the kinetic mixing parameter. The interaction is equivalent to one between the EM field $A_{\mu}$ and an effective current $J^{\mu}_{\rm eff}$. Applying the Euler-Lagrange equations with respect to $A^{\mu}$ to $\mathcal{L}\supset-\frac{1}{4}F^{\mu\nu}F_{\mu\nu}-J^{\mu}_{\rm eff}A_{\mu}$, we obtain the equation of motion $\partial_{\nu}F^{\mu\nu}=J^{\mu}_{\rm eff}$. Using the equation of motion for massive vector fields $\partial_{\mu}{A^{\prime}}^{\mu}=0$, taking the non-relativistic limit, one can see ${A^{\prime}}^{0}=0$. Then, the dark photon field can be written as higgins2024maglev
@@
Axion-Photon Coupling.– Axions can also couple to photons via the following operator which can also be written as an interaction between the EM field and an effective current kalia2024ultralight
@@ Appendix B Appendix B: MRFM Details
We first expand on inductive vs. mechanical detection. In conventional inductive detection, the precession of a sample’s magnetization induces a current in a pickup coil, which in turn generates a magnetic field that reacts back on the sample. This exchange of energy resulting in back-action noise, historically known as radiation damping, effectively damps the magnetization precession within the characteristic time $T_{R}=(2\pi\eta Q_{\rm coil}\gamma\mu_{0}M_{0})^{-1}$, where $\eta$ is the filling factor (volume of sample per volume of coil ($V_{s}/V_{c}$)) and $Q_{\rm coil}$ is the coil’s quality factor muller2007nuclear; ruoso2016quax. MRFM overcomes this limitation by interacting with the sample through a magnetic force gradient (a one-way force detection) rather than a reciprocal electromagnetic induction (a two-way energy exchange). This allows for a less perturbative and therefore more sensitive measurement with a better signal-to-noise ratio sidles1993signal.
While various modulation schemes enable detection of longitudinal magnetization, transverse magnetization detection in MRFM is challenging. This is due to the frequency mismatch between typical Larmor and mechanical frequencies. Even in the absence of this mismatch, spin control is complicated due to the requirement of a strong radio frequency field and non-applicability of the rotating wave approximation fischer2019spin.
Micromagnet.– Here we expand on the theoretical treatment of the micromagnet. An integral part of an MRFM setup is the micromagnet that generates an inhomogeneous magnetic field with a spatial gradient that couples the mechanical resonator to the sample magnetic moments. In our setup, the micromagnet is approximated as a micro-sphere with radius $R_{s}$ and has saturation magnetization $M_{0}$, uniformly magnetized along $\hat{z}$. The criterion for the selection of the radius $R_{s}$ is its ratio with the distance between the center of the micromagnet and the surface of the sample ($d$), such that $R_{s}/d\ll 1$, and hence the magnet creates a dipolar magnetic field with strength $B(r)=\frac{4\pi}{3}M_{0}(R_{s}/r)^{3}$ at radial distance $r$ from its center suter2002probe; hammel2007magnetic; sidles91. This gives the following field gradient
@@
For the material of the micromagnet, we considered Fe-Co because of its high saturation magnetization $M_{0}\sim 2.4\quad A/m$ xue2011geometry; liu2008magnetic.
Micromagnet Loaded Mechanical System.– Here we discuss how to calculate the resonance frequency of a micromagnet loaded mechanical resonator. For example, consider a $V_{m}=\rm 5~mm\times 5~mm\times 500~nm$ volume SiN membrane ($\rho_{m}=3.17~\rm g/cm^{3}$) with spring constant $k=60~\rm N/m$ loaded with a micromagnet of mass $m_{\rm magnet}=36~\rm\mu g$. Then the mechanical resonator frequency can be calculated using $\omega_{m}=\sqrt{k/m_{\rm eff}}$, where the effective mass is approximately $m_{\rm eff}\approx m_{\rm magnet}+\frac{m_{\rm mech}}{4}$. Using the above numbers, the magnet-loaded membrane frequency is $\omega_{m}\approx 2\pi\times 5.8~\rm kHz$. For cantilever, we use the magnetically loaded resonance frequency $\omega_{m}=2\pi\times 1.1~\rm kHz$, which matches that of wago1998paramagnetic. In general, the resonance frequency of the loaded mechanical system will be estimated numerically.
Resonance Volume.– The micromagnet and the bias field $B_{0}$ define the region of space where magnetic resonance can be induced. For conditions $B_{0}\neq B_{\rm rf}/\gamma$, the sensitive slice is a thin shell with spatial thickness $x_{sl}\approx\frac{\delta B}{\nabla B}$, where $\delta B\propto\frac{1}{T_{2}}$ is the intrinsic resonance linewidth of the sample. However, this concept breaks down when $B_{0}=B_{\rm rf}/\gamma$, where the influence of the gradient becomes small across most of the sample, hence known as zero-probe-field resonance (ZPFR). In this regime, the sensitive region is no longer a slice and expands to cover nearly the entire sample. Thus, each spin experiences a total field close to $B_{0}$, and this results in a strong signal despite the weak local field suter2002probe.
Changes in longitudinal magnetization.– The evolution of an ensemble of spins in a magnetized sample due to external magnetic fields (magnetizing field $B_{0}\hat{z}$, and a driving radio-frequency field $B_{\rm rf}$ in the x-y plane) is described by Bloch equations abragam1961principles:
@@
For low d.c. fields, comparable in magnitude to the r.f. field, one instead considers the modified Bloch equations $\frac{d\vec{m}}{dt}=\gamma\vec{m}\times\vec{B}-\frac{m_{z}-\chi_{0}(B_{0}+B_{z})}{T_{1}}\hat{z}-\frac{m_{x}\hat{x}+m_{y}\hat{y}}{T_{2}}$ abragam1961principles. The longitudinal steady-state solution is
@@
where $\frac{\chi_{0}}{\mu_{0}}B_{0}=m_{0}$, $B_{0}\gamma=\omega_{L}$, $\Delta\omega=\omega_{L}-\omega_{\rm rf}$. At low driving field (as defined by $\gamma^{2}B_{\rm rf}^{2}T_{1}T_{2}\ll 1$), the spins excited at resonance ($\omega=\omega_{L}=\omega_{\rm rf}$) vary the longitudinal magnetization to $m_{0}+\Delta m_{z}$, where $\Delta m_{z}=m_{0}\gamma^{2}B_{\rm rf}^{2}T_{1}T_{2}$ wago1997magnetic.
@@
Magnetic-Dipole radiation damping.- For a magnetic sample in free space with a high Larmor frequency and large magnetization $M_{0}$, the dominant loss mechanism is the magnetic dipole emission. This is an intrinsic quantum phenomenon where the precessing magnetization spontaneously emits electromagnetic radiation characterized by the relaxation time $\tau_{r}=\frac{\lambda_{L}^{3}}{\gamma\mu_{0}M_{0}V_{s}}$, where $\lambda_{L}=c/\nu_{L}$ is the Larmor wavelength abragam1961principles; crescini2019towards; ruoso2016quax. To avoid faster than spin-spin relaxation, the spherical YIG sample with Larmor frequency $1~\rm GHz$ ($10~\rm GHz$) must be limited in size to a maximum radius of around $9~\rm mm$ ($0.6~\rm mm$). Consequently, we restrict our investigation to frequencies around a $\rm GHz$ to avoid this coherence loss channel. Alternatively, strategies like embedding the sample in a resonant cavity (as done by QUAX) may be employed to manage the phase space of the emitted radiation crescini2019towards; ruoso2016quax; crescini2018operation.
@@
Figure 4: The magnetic field gradient produced by the resonator-attached micromagnet (shown as a red dot at the origin) is plotted as a contour against distance. The yellow region denotes the sensitive slice that is determined by the magnetic resonance condition $B_{0}+B_{z}=\frac{\omega_{L}(\vec{r})}{\gamma}$ and the linewidth $1/T_{2}$. The red circles depict a $3\times 3$ grid of YIG spheres.
@@ Appendix C Appendix C: Force Equivalent Noise
The transfer function between force and position is obtained by solving the equations of motion for the system in the frequency domain aspelmeyer2014cavity
@@
saulson1990thermal, which are the imprecision, backaction, and thermal noise respectively.
Here, $G=\omega_{c}/L$ is the optomechanical coupling aspelmeyer2014cavity with cavity frequency $\omega_{c}$ and cantilever length $L$, $\kappa=\frac{\omega_{\rm FSR}}{F}=\frac{c}{2LF}$ with finesse $F$ is the cavity decay rate, and
@@
The quarter-root scaling with integration time is due to the fact that the coherence time of the signal is much less than the integration time. Through a method of averaging the power spectral densities (PSDs) of several short-time measurements of length equal to 1 coherence time (known as Bartlett’s method proakis2006), we arrive at the quarter-root scaling manley2023searching.

### v2

@@ Appendix A Appendix A: Dark Matter Couplings
Dark Photon Kinetic Mixing.– The dark photon ${A^{\prime}}^{\mu}$ can kinetically mix with with the Standard Model photon $A^{\mu}$, leading to the following interaction operator in the interaction basis [40, 62]
@@
where $\epsilon$ is the kinetic mixing parameter. The interaction is equivalent to one between the EM field $A_{\mu}$ and an effective current $J^{\mu}_{\rm eff}$. Applying the Euler-Lagrange equations with respect to $A^{\mu}$ to $\mathcal{L}\supset-\frac{1}{4}F^{\mu\nu}F_{\mu\nu}-J^{\mu}_{\rm eff}A_{\mu}$, we obtain the equation of motion $\partial_{\nu}F^{\mu\nu}=J^{\mu}_{\rm eff}$. Using the equation of motion for massive vector fields $\partial_{\mu}{A^{\prime}}^{\mu}=0$, taking the non-relativistic limit, one can see ${A^{\prime}}^{0}=0$. Then, the dark photon field can be written as [53]
@@
Axion-Photon Coupling.– Axions can also couple to photons via the following operator which can also be written as an interaction between the EM field and an effective current [62]
@@ Appendix B Appendix B: MRFM Details
We first expand on inductive vs. mechanical detection. In conventional inductive detection, the precession of a sample’s magnetization induces a current in a pickup coil, which in turn generates a magnetic field that reacts back on the sample. This exchange of energy resulting in back-action noise, historically known as radiation damping, effectively damps the magnetization precession within the characteristic time $T_{R}=(2\pi\eta Q_{\rm coil}\gamma\mu_{0}M_{0})^{-1}$, where $\eta$ is the filling factor (volume of sample per volume of coil ($V_{s}/V_{c}$)) and $Q_{\rm coil}$ is the coil’s quality factor [74, 82]. MRFM overcomes this limitation by interacting with the sample through a magnetic force gradient (a one-way force detection) rather than a reciprocal electromagnetic induction (a two-way energy exchange). This allows for a less perturbative and therefore more sensitive measurement with a better signal-to-noise ratio [87].
While various modulation schemes enable detection of longitudinal magnetization, transverse magnetization detection in MRFM is challenging. This is due to the frequency mismatch between typical Larmor and mechanical frequencies. Even in the absence of this mismatch, spin control is complicated due to the requirement of a strong radio frequency field and non-applicability of the rotating wave approximation [42].
Micromagnet.– Here we expand on the theoretical treatment of the micromagnet. An integral part of an MRFM setup is the micromagnet that generates an inhomogeneous magnetic field with a spatial gradient that couples the mechanical resonator to the sample magnetic moments. In our setup, the micromagnet is approximated as a micro-sphere with radius $R_{s}$ and has saturation magnetization $M_{0}$, uniformly magnetized along $\hat{z}$. The criterion for the selection of the radius $R_{s}$ is its ratio with the distance between the center of the micromagnet and the surface of the sample ($d$), such that $R_{s}/d\ll 1$, and hence the magnet creates a dipolar magnetic field with strength $B(r)=\frac{4\pi}{3}M_{0}(R_{s}/r)^{3}$ at radial distance $r$ from its center [90, 52, 85]. This gives the following field gradient
@@
For the material of the micromagnet, we considered Fe-Co because of its high saturation magnetization $\mu_{0}M_{0}\sim 2.4~\rm T$ [104, 67].
Micromagnet Loaded Mechanical System.– Here we discuss how to calculate the resonance frequency of a micromagnet loaded mechanical resonator. For example, consider a $V_{m}=\rm 5~mm\times 5~mm\times 500~nm$ volume SiN membrane ($\rho_{m}=3.17~\rm g/cm^{3}$) with spring constant $k=60~\rm N/m$ loaded with a micromagnet of mass $m_{\rm magnet}=36~\rm\mu g$. Then the mechanical resonator frequency can be calculated using $\omega_{m}=\sqrt{k/m_{\rm eff}}$, where the effective mass is approximately $m_{\rm eff}\approx m_{\rm magnet}+\frac{m_{\rm mech}}{4}$. Using the above numbers, the magnet-loaded membrane frequency is $\omega_{m}\approx 2\pi\times 5.8~\rm kHz$. For cantilever, we use the magnetically loaded resonance frequency $\omega_{m}=2\pi\times 1.1~\rm kHz$, which matches that of [95]. In general, the resonance frequency of the loaded mechanical system will be estimated numerically.
Resonance Volume.– The micromagnet and the bias field $B_{0}$ define the region of space where magnetic resonance can be induced. For conditions $B_{0}\neq B_{\rm rf}/\gamma$, the sensitive slice is a thin shell with spatial thickness $x_{sl}\approx\frac{\delta B}{\nabla B}$, where $\delta B\propto\frac{1}{T_{2}}$ is the intrinsic resonance linewidth of the sample. However, this concept breaks down when $B_{0}=B_{\rm rf}/\gamma$, where the influence of the gradient becomes small across most of the sample, hence known as zero-probe-field resonance. In this regime, the sensitive region is no longer a slice and expands to cover nearly the entire sample. Thus, each spin experiences a total field close to $B_{0}$, and this results in a strong signal despite the weak local field [90].
Changes in longitudinal magnetization.– The evolution of an ensemble of spins in a magnetized sample due to external magnetic fields (magnetizing field $B_{0}\hat{z}$, and a driving radio-frequency field $B_{\rm rf}$ in the x-y plane) is described by Bloch equations [3]:
@@
For low d.c. fields, comparable in magnitude to the r.f. field, one instead considers the modified Bloch equations $\frac{d\vec{m}}{dt}=\gamma\vec{m}\times\vec{B}-\frac{m_{z}-\chi_{0}(B_{0}+B_{z})}{T_{1}}\hat{z}-\frac{m_{x}\hat{x}+m_{y}\hat{y}}{T_{2}}$ [3]. The longitudinal steady-state solution is
@@
where $\frac{\chi_{0}}{\mu_{0}}B_{0}=m_{0}$, $B_{0}\gamma=\omega_{L}$, $\Delta\omega=\omega_{L}-\omega_{\rm rf}$. At low driving field (as defined by $\gamma^{2}B_{\rm rf}^{2}T_{1}T_{2}\ll 1$), the spins excited at resonance ($\omega=\omega_{L}=\omega_{\rm rf}$) vary the longitudinal magnetization to $m_{0}+\Delta m_{z}$, where $\Delta m_{z}=m_{0}\gamma^{2}B_{\rm rf}^{2}T_{1}T_{2}$ [96].
@@
Magnetic-dipole radiation damping.- For a magnetic sample in free space with a Larmor frequency and large magnetization $M_{0}$, the dominant loss mechanism is the magnetic dipole emission. This phenomenon can be understood classically as lost power in the form of emitted photons driven by the magnetic dipole, and is characterized by a relaxation time $T_{r}=\frac{c^{3}}{\omega^{3}}\frac{6\pi}{\mu_{0}\gamma M_{0}V_{s}}$ [51]. This gives a restriction on how large of a sample volume $V_{s}$ one can use. Alternatively, strategies like embedding the sample in a resonant cavity (as done by QUAX) may be employed to avoid this loss mechanism [37, 82, 36].
@@
Figure 4: The magnetic field gradient produced by the resonator-attached micromagnet (shown as a red dot at the origin) is plotted as a contour against distance. The yellow region denotes the sensitive slice that is determined by the magnetic resonance condition $B_{0}+B_{z}=\frac{\omega_{L}(\vec{r})}{\gamma}$ and the linewidth $1/T_{2}$. The red rectangle depicts the YIG sample.
@@ Appendix C Appendix C: Force Equivalent Noise
The transfer function between force and position is obtained by solving the equations of motion for the system in the frequency domain [11]
@@
[83], which are the imprecision, backaction, and thermal noise respectively.
Here, $G=\omega_{c}/L$ is the optomechanical coupling [11] with cavity frequency $\omega_{c}$ and cantilever length $L$, $\kappa=\frac{\omega_{\rm FSR}}{F}=\frac{c}{2LF}$ with finesse $F$ is the cavity decay rate, and
@@
The quarter-root scaling with integration time is due to the fact that the coherence time of the signal is much less than the integration time. Through a method of averaging the power spectral densities (PSDs) of several short-time measurements of length equal to 1 coherence time (known as Bartlett’s method [79]), we arrive at the quarter-root scaling [68].